OpenStudy (anonymous):

Write in simplest fraction form A) (0,4)^-1 B) (0,8)^-2

OpenStudy (anonymous):

Negative power is "to put in the denominator"

OpenStudy (anonymous):

That means \[a^{-1} = \frac{ 1 }{ a }\]

OpenStudy (anonymous):

How much is 1/0.4 @sarmandi ?

OpenStudy (anonymous):

2,5

OpenStudy (anonymous):

Well now u have answer to A

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

About B:\[a^{-2} = \frac{ 1 }{ a^2 }\]

OpenStudy (shane_b):

I'm confused as to the notation here...what do they mean by (0,4)?

OpenStudy (anonymous):

@sarmandi is this Q. from complex analysis ?

OpenStudy (anonymous):

\[(0,4)^{-1}\]

OpenStudy (anonymous):

Check your meanining of (0,4)^{-1}

OpenStudy (shane_b):

Maybe it was wrong because of how you put it in the system...it says to write it as a fraction:\[(0,4)^{-1}=\frac{1}{0,4}\]\[(0,8)^{-2}=\frac{1}{0,8^2}\]That's assuming the notation used in the question.